OpenStudy (anonymous):

What is 19.8% of 7500 meters? (Round too the nearest tenth)

OpenStudy (anonymous):

To figure this out, you would multiply 7500 by .198. You can insert that into a calculator or even type it into google like 7500*.198.

OpenStudy (anonymous):

% --> 1/100 \[{19.8\%}\]\[=\frac{19.8}{100}\] so, 19.8 of 7500 = \[=\frac{19.8*7500}{100}\]

OpenStudy (anonymous):

now calculate

OpenStudy (anonymous):

@MauriceLong did you understand that?